### 03 Mar 2026 v2

@@ Abstract
Based on entropy and symmetrical uncertainty (SU), we define a metric for categorical random variables and show that this metric can be promoted into an appropriate quotient space of categorical random variables. Moreover, we also show that there is a natural commutative monoid structure in the same quotient space, which is compatible with the topology induced by the metric, in the sense that the monoid operation is continuous.
@@ 1 Introduction
Symmetric Uncertainty has been introduced in [6], and it is defined as
@@
where $X,Y$ are categorical random variables, (i.e.: random variables taking values in a finite set, not necessarily numerical), $H(\cdot)$ is the Shannon entropy [10], and $H(\cdot,\cdot)$ is the joint Shannon entropy.
Generalization of SU for the multivariate case, as well as its properties, and, in particular, its potential to define a distance metric among categorical random variables have been explored in [5, 4, 12, 11]. Indeed, [4] bring us to realize that entropic correlation is an example of similarity.
In this work, we further explore some properties of $SU$ in two directions. On one hand we explore some properties of the topology introduced by the metric defined in terms of $SU$ on a given space of categorical random variables. On the other hand, we explore some algebraic properties of a joint operation defined on the same space where the metric is defined. We conclude that this operation endows the metric space with a commutative monoid structure, and that it is compatible with the metric topology, in the sense that it is a continuous map.
@@ 2 Background: Entropic Correlation and Symmetric Uncertainty
In this section we recall the fundamental concepts involved in the definition of $SU$, as well as its main properties which we are going to need.
Let us consider a fixed probability space $(D,\Omega_{D},P)$, and let $X\colon D\to\Sigma_{X}$ be a categorical (non-numeric, nominal or cualitative) random variable with possible values on the finite set $\Sigma_{X}=\{x_{1},\ldots,x_{k}\}$ and probability mass function $P_{X}\colon\Sigma_{X}\to\mathbb{R}$ defined as $P_{X}(x_{i}):=P(X=x_{i}):=P(X^{-1}(x_{i}))$. The entropy $H$ of the variable $X$ is a measure of the uncertainty in predicting the value of $X$, or size of the uncertainty on $X$, and is defined as [9]:
@@
Mutual Information. Using entropy and conditional entropy introduced above, the mutual information  [^1] In certain literature mutual information is denoted as $I(X;Y)$; see [3]. $MI(X|Y)$ is introduced [7]. Mutual Information measures the reduction in uncertainty about the value of $X$ when the value of $Y$ is known, and is defined as:
MI(X|Y):=H(X)H(X|Y).MI(X|Y):=H(X)-H(X|Y). (5)
@@
  • if $X$ and $Y$ are independent then in equation (5) $H(X|Y)=H(X)$, and hence $MI(X|Y)$ = 0. That is, under independence, knowing $Y$ gives no information about the value of $X$; and
  • if $X$ and $Y$ are fully correlated then $H(X|Y)=0$ and hence $MI(X|Y)=H(X)$ applying equation (5).
Mutual information taken as a measure of correlation, makes it possible to compare correlations with one another. For example, for any random variables $X$, $Z$ and $Y$, $MI(X|Y)>MI(Z|Y)$ means that when knowing the value of $Y$ the reduction in uncertainty about $X$ is greater than the reduction in uncertainty about $Z$, which is conducive to stating that $X$ is more correlated to $Y$ than it is $Z$.
@@
MI(X|Y)=H(X)+H(Y)H(X,Y)MI(X|Y)=H(X)+H(Y)-H(X,Y) (6)
and since $H(X,Y)$ is symmetric by direct substitution of $X,Y$ into equation (1), it results that $MI(X|Y)$ is symmetric as well, a quite convenient property for a paired measure. A formal proof can be found, for example, as Theorem 1 in [3].
Symmetric Uncertainty. On the other hand, like with other informational association measures, MI tends to increase its value when the number of categories in $X$ and/or $Y$ increases, that is, it is biased towards high cardinality features [8]. Therefore, seeking to compensate for such bias, MI has to be normalized by the sum of entropies of the features. This measure, called Symmetric Uncertainty [6] is expressed as:
$SU(X,Y):=2\left[\frac{MI(X|Y)}{H(X)+H(Y)}\right]$ (7)
@@
Revisiting the two bullet points in the previous subsection, note now that:
  • if $X$ and $Y$ are independent then, from $MI(X|Y)$ = 0, it follows that $SU(X,Y)=0$; and
  • if $X$ and $Y$ are completely correlated then from equation (5), it is obtained $MI(X|Y)=H(X)$. By symmetry of MI, $MI(Y|X)=H(Y)$ and so $SU(X,Y)=1$.
@@
Notice that, taking ${\cal Z}={\cal D}$ in item $a)$ we get, for any two partition $\cal X$} and $\cal Y$}:
$H({\cal X\lor{\cal Y}})=H({\cal X})+H({\cal Y|X)}\implies H({\cal X})\leq H({\cal X}\lor{\cal Y})$ (11)
@@
is a categorical random variable defined on $D$, then it defines a partition $\cal X$} of $D$ given by the inverse images of $x_{i}$, i.e.:
@@ 3 Topological structure of Categorical Variables
In this section we introduce a topology into a set of categorical variables. We will do so by defining a distance metric based on the SU measure. To motivate the definition, we will first explore some examples in order to get an idea of how the SU measure can give us a way to compare categorical variables.
@@ 3.1 Measuring similarity on a dataset: between observations and between variables
A set of data, or dataset, consists of $m$ observations taken across $n$ variables, or put another way, of $n$ random variables $X_{j}$ evaluated on $m$ subjects $q_{i}$. Let us set $x_{ij}=X_{j}(q_{i})$, so we can arrange the observed values as shown in Table 1.
$X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ $X_{5}$ ... $X_{n}$
@@
Looking row-wise at the dataset, any two observations (rows) can be compared to determine their similarity. If two rows are equal throughout all their values, we say that they are in full similarity or equivalently, we say that the distance between these observations is zero. If not all the values on the two rows are equal, several techniques are available to compute distances in this $n$-dimensional space, where a mix of data types is not unusual; see for instance Suárez-Díaz et al [13].
@@
Is it possible to find similarities between columns? Since each column measures a specific characteristic, and perhaps has been recorded in its own unit of measurement, we can rarely expect to find column similarity in the same way that we find similar rows. However, here one can think of similarity in a slightly different way. It is known that columns may relate to each other via correlations. Correlation does not necessarily imply causality, but it is reasonable to view high correlation between two variables $X_{i}$ and $X_{j}$ as an indicator of being linked to the same phenomenon. Because of this, we may think that if the correlation between $X_{i}$ and $X_{j}$ is high, then the two variables are “close of each other”; and conversely, low correlation between $X_{i}$ and $X_{j}$ sets these variables apart.
@@
For categorical variables, the Symmetrical Uncertainty measure of correlation computes this kind of similarity by comparing the joint entropy of the pair of variables with their individual entropies. So the answer to the above questions is: Yes, similarity expressed as correlation between two columns is computable as the SU. This similarity is based on variables belonging to a group of intertwined characteristics. The next example illustrates how two columns may be viewed as similar.
Example 1.
Students applying for certain internship got their scores recorded on 5 traits: Neatness, Punctuality, IQ, Attention Type, Creativity, and GotHired.
@@ 20 R I L H D N
Table 2 presents data example for the personality traits of students applying for an internship. Computing probabilities from the observed frrequency, Table 3 shows the calculation of $SU$ for the variables Neatness and Hired based on the three terms in Equation (8).
@@
The full set of resulting $SU$ values between each of the variables and the Hired response are:
@@
Out of many possible predictive models, a model to predict whether a student will get hired based on a single predictor variable is perhaps the simplest one. Of all the 5 feature vs class pairs in the data, the creativity-gotHired pair has highest SU value. Let us analyze this a bit more.
Suppose we need to group our 20 cases by each category of Creativity and by each category of GotHired. When we do this, Creativity and GotHired produce about the same groupings, as seen in Table 4: creativity D is associated with GotHired=Y, and creativities S or I associate with GotHired=N.
Table 4: Counts for Creativity and Getting Hired
@@ D S I
Thus creativity is akin to getting hired, or Creativity is similar to GotHired in the correlation sense. From the point of view of a 2-way model, if we know that the value of creativity is D we can almost predict that the value of GotHired will be Y, while for creativities S or I the candidate probably won’t be hired.
The same kind of analyses could be done for the other four resulting SU values, detecting similarities with GotHired in accordance with each SU value.
Note that if we replace GotHired by its negative NotHired, all the $SU$ values would remain the same as only the probabilities of the labels (not labels themselves) participate in $SU$ computations.
@@ 3.2 Entropic equivalence and the metric structure
In many situations, given a set $\Omega$, it is useful to have a function $f\colon\Omega\times\Omega\to\mathbb{R}$ which tell us how similar or dissimilar are any given pair of elements of $\Omega$. One such a function is the well known distance metric, which is any function $d\colon\Omega\times\Omega\to\mathbb{R}$ satisfying the following conditions:
@@
Less standardized is the concept of a similarity metric. Here we recall the definition given by Chen et al in [2].
@@
In [2], the authors also gave conditions for a similarity metric to yield a distance metric. More concretely, they established a precise relationship between normalized distances and similarity metrics. Here, normalized means that both the distance and the similarity map into the interval $0,1]\subset\mathbb{R}$. For the reader convenience, let us recall their result here.
Lemma 3.2 (Corollary 1 in [2]).
@@
Let $\cal C$ be a set of categorical random variables defined on a sample space $D$. If $X\colon D\to\Sigma_{X}$ and $Y\colon D\to\Sigma_{Y}$ are two elements in $\cal C$}, we say that they are indiscernible if there is a bijection $h\colon\Sigma_{X}\to\Sigma_{Y}$ and $Y=h\circ X$ almost everywhere.
@@
Example 2.
Example of indiscernibles. Consider variable $X$ with values {1, 2, 3} and variable $Y$ with values {A, B, C}. Table 5 shows a sample of 10 individuals. Considering the probability given by the observed frequency, we can see that the histogram of $X$ is 0.5, 0.1, 0.4, while that of $Y$ is 0.4, 0.1, 0.5. Here, we can define the bijection $h\colon\{1,2,3\}\to\{A,B,C\}$ given by $h(1)=C;h(2)=B;h(3)=A$, and then the partition defined on the domain by $Y$ and $h\circ X$ are the same.
Table 5: Example of Indiscernibles.
@@
    Positivity and normality: $SU(X,Y)\in[0,1]$.
Proof.
@@
Condition 1 is straightforward by substitution into equation (7). Likewise, for the reflexivity condition 2, using equation (7) we get the self-similarity $SU(X,X)=1\geq 0$.
The self-similarity condition 3 establishes that $SU(X,X)$ is greater than or equal to any other value of SU(X, Y). This is indeed true as we have just seen that $SU(X,X)=1$ which is the maximum value achievable by the measure.
Equivalent to the triangle inequality in a distance metric, the Triangle Inequality (Condition 4) states that the similarity between $X$ and $Z$ through $Y$ is no greater than the direct similarity between $X$ and $Z$ plus the self similarity of $Y$. Let us start by rewriting inequalities for the two-variable cases, using entropic ratio notation (see Definition in Lemma 2.1). For the variables $X$, $Y$, and $Z$, we get:
@@
b. From here on use only the left and the middle members. Transpose two terms to the left, getting
@@
Condition 5 results from applying $SU(X,X)$ = 1 once again, and noting that $X=Y$ is understood as equality in entropy.
Finally, notice that Condition 6 was already established in Lemma 2.1. This finalizes the proof. ∎
@@
Given a space $\cal C$ of categorical random variables, let $\overline{\cal C}$ be its coset defined by indiscernibility. The quantity 1SU(X,Y)1-SU(X,Y) is a normalized distance metric in $\overline{\cal C}$.
Let us notice that, since we are working with categorical variables, it may seem that the metric topology just introduced is discrete. However, this is not the case, as we show in the following theorem.

### 30 Mar 2026 v4

@@ Abstract
Based on entropy and symmetric uncertainty (SU), we define a metric for categorical random variables and show that this metric can be promoted into an appropriate quotient space of categorical random variables. Moreover, we show that there is a natural commutative monoid structure in the same quotient space, which is compatible with the topology induced by the metric, in the sense that the monoid operation is continuous.
keywords:
@@ 1 Introduction
Symmetric Uncertainty (SU) has been introduced in [6], and it is defined as:
@@
where $X,Y$ are categorical random variables (i.e., random variables taking values in a finite set, not necessarily numerical), $H(\cdot)$ is the Shannon entropy [10], and $H(\cdot,\cdot)$ is the joint Shannon entropy.
The generalization of SU for the multivariate case, as well as its properties and, in particular, its potential to define a distance metric among categorical random variables have been explored in [4, 5, 12, 11]. Indeed, [5] shows that entropic correlation is an example of similarity.
In this work, we further explore some properties of $SU$ in two directions. On the one hand, we explore some properties of the topology introduced by the metric defined in terms of $SU$ on a given space of categorical random variables. On the other hand, we explore some algebraic properties of a joint operation defined on the same space where the metric is defined. We conclude that this operation endows the metric space with a commutative monoid structure, and that it is compatible with the metric topology, in the sense that it is a continuous map.
@@ 2 Background: Entropic Correlation and Symmetric Uncertainty
In this section, we recall the fundamental concepts involved in the definition of $SU$, as well as its main properties, which we will need.
Let us consider a fixed probability space $(D,\Omega_{D},P)$, and let $X\colon D\to\Sigma_{X}$ be a categorical (non-numeric, nominal or qualitative) random variable with possible values in the finite set $\Sigma_{X}=\{x_{1},\ldots,x_{k}\}$ and probability mass function $P_{X}\colon\Sigma_{X}\to\mathbb{R}$ defined as $P_{X}(x_{i}):=P(X=x_{i}):=P(X^{-1}(x_{i}))$. The entropy $H$ of the variable $X$ is a measure of the uncertainty in predicting the value of $X$, or size of the uncertainty on $X$, and is defined as [9]:
@@
Mutual Information. Using entropy and conditional entropy introduced above, the mutual information  [^1] In certain literature mutual information is denoted as $I(X;Y)$; see [3]. $MI(X;Y)$ is introduced [8]. Mutual information measures the reduction in uncertainty about the value of $X$ when the value of $Y$ is known, and is defined as:
MI(X;Y):=H(X)H(X|Y).MI(X;Y):=H(X)-H(X|Y). (5)
@@
  • if $X$ and $Y$ are independent then in equation (5) $H(X|Y)=H(X)$, and hence $MI(X;Y)=0$. That is, under independence, knowing $Y$ gives no information about the value of $X$; and
  • if $X$ and $Y$ are fully correlated then $H(X|Y)=0$ and hence $MI(X;Y)=H(X)$ applying equation (5).
Mutual information, taken as a measure of correlation, makes it possible to compare correlations with one another. For example, for any random variables $X$, $Z$ and $Y$, $MI(X;Y)>MI(Z;Y)$ means that, when the value of $Y$ is known, the reduction in uncertainty about $X$ is greater than the reduction in uncertainty about $Z$, which suggests that $X$ is more correlated with $Y$ than with $Z$.
@@
MI(X;Y)=H(X)+H(Y)H(X,Y),MI(X;Y)=H(X)+H(Y)-H(X,Y), (6)
and since $H(X,Y)$ is symmetric by direct substitution of $X,Y$ into equation (1), it results that $MI(X;Y)$ is symmetric as well, a quite convenient property for a paired measure. A formal proof can be found, for example, as Theorem 1 in [3].
Symmetric Uncertainty. On the other hand, as with other informational association measures, MI tends to increase its value when the number of categories in $X$ and/or $Y$ increases, that is, it is biased towards high cardinality features [7]. Therefore, seeking to compensate for such bias, MI has to be normalized by the sum of entropies of the features. This measure, called Symmetric Uncertainty [6] is expressed as:
$SU(X,Y):=2\left[\frac{MI(X;Y)}{H(X)+H(Y)}\right]$ (7)
@@
Revisiting the two bullet points in the previous subsection, note that:
  • If $X$ and $Y$ are independent, then from $MI(X;Y)=0$ it follows that $SU(X,Y)=0$.
  • If $X$ and $Y$ are completely correlated, then from (5) we obtain
    $MI(X;Y)=H(X).$
    By symmetry of mutual information, $MI(Y;X)=H(Y)$, and therefore
    $SU(X,Y)=1.$
@@
Notice that, taking ${\cal Z}={\cal D}$ in item $a)$ we get, for any two partitions $\cal X$} and $\cal Y$}:
$H({\cal X\lor{\cal Y}})=H({\cal X})+H({\cal Y|X})\implies H({\cal X})\leq H({\cal X}\lor{\cal Y}).$ (11)
@@
is a categorical random variable defined on $D$, then it defines a partition $\cal X$} of $D$ given by the inverse images of $x_{i}$, i.e.,
@@ 3 Topological structure of Categorical Variables
In this section we introduce a topology on a set of categorical variables. We will do so by defining a distance metric based on the SU measure. To motivate the definition, we will first explore some examples in order to get an idea of how the SU measure can give us a way to compare categorical variables.
@@ 3.1 Measuring similarity on a dataset: between observations and between variables
A set of data, or dataset, consists of $m$ observations taken across $n$ variables, or equivalently, $n$ random variables $X_{j}$ evaluated on $m$ subjects $q_{i}$. Let us set $x_{ij}=X_{j}(q_{i})$, so we can arrange the observed values as shown in Table 1.
$X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ $X_{5}$ $X_{n}$
@@
Looking row-wise at the dataset, any two observations (rows) can be compared to determine their similarity. If two rows are equal throughout all their values, we say that they are in full similarity or equivalently, we say that the distance between these observations is zero. If not all the values on the two rows are equal, several techniques are available to compute distances in this $n$-dimensional space, where a mix of data types is not unusual; see for instance Suárez-Díaz et al. [13].
@@
Is it possible to find similarities between columns? Since each column measures a specific characteristic, and perhaps has been recorded in its own unit of measurement, we rarely expect to find column similarity in the same way that we find similar rows. However, here one can think of similarity in a slightly different way. It is known that columns may relate to each other via correlations. Correlation does not necessarily imply causality, but it is reasonable to view high correlation between two variables $X_{i}$ and $X_{j}$ as an indicator of being linked to the same phenomenon. Because of this, we may think that if the correlation between $X_{i}$ and $X_{j}$ is high, then the two variables are “close to each other”; and conversely, low correlation between $X_{i}$ and $X_{j}$ sets these variables apart.
@@
For categorical variables, the Symmetric Uncertainty measure of correlation computes this kind of similarity by comparing the joint entropy of the pair of variables with their individual entropies. So the answer to the above questions is: Yes, similarity expressed as correlation between two columns is computable as the SU. This similarity is based on variables belonging to a group of intertwined characteristics. The next example illustrates how two columns may be viewed as similar.
Example 3.1.
Students applying for a certain internship had their scores recorded on 5 traits: Neatness, Punctuality, IQ, Attention Type, Creativity, along with a response variable GotHired.
@@ 20 R I L H D N
Table 2 presents an example dataset for the personality traits of students applying for an internship. Computing probabilities from the observed frequency, Table 3 shows the calculation of $SU$ for the variables Neatness and Hired based on the three terms in Equation (8).
@@
The full set of resulting $SU$ values between each of the variables and the Hired response is:
@@
Out of many possible predictive models, a model to predict whether a student gets hired based on a single predictor variable is perhaps the simplest one. Of all the 5 feature vs class pairs in the data, the creativity-gotHired pair has the highest SU value. Let us analyze this a bit more.
Suppose we need to group our 20 cases by each category of Creativity and by each category of GotHired. When we do this, Creativity and GotHired produce about the same groupings, as seen in Table 4: creativity D is associated with GotHired=Y, and creativity levels S or I are associated with GotHired=N.
@@ D S I
Table 4: Counts for Creativity and Getting Hired
Thus, creativity is akin to getting hired, or Creativity is similar to GotHired in the correlation sense. From the point of view of a two-way model, if we know that the value of creativity is D, we can almost predict that the value of GotHired will be Y, while for creativities S or I the candidate probably will not be hired.
The same kind of analysis could be done for the other four resulting SU values, detecting similarities with GotHired in accordance with each SU value.
Note that, if we replace GotHired by its negative NotHired, all the $SU$ values would remain the same as only the probabilities of the labels (not labels themselves) participate in $SU$ computations.
@@ 3.2 Entropic equivalence and the metric structure
In many situations, given a set $\Omega$, it is useful to have a function $f\colon\Omega\times\Omega\to\mathbb{R}$ which tells us how similar or dissimilar any given pair of elements of $\Omega$ is. One such function is the well-known distance metric, which is any function $d\colon\Omega\times\Omega\to\mathbb{R}$ satisfying the following conditions:
@@
Less standardized is the concept of a similarity metric. Here, we recall the definition given by Chen et al. in [2].
@@
In [2], the authors also gave conditions for a similarity metric to yield a distance metric. More concretely, they established a precise relationship between normalized distances and similarity metrics. Here, normalized means that both the distance and the similarity map into the interval $0,1]\subset\mathbb{R}$. For the reader’s convenience, let us recall their result here.
Lemma 3.3 ([2, Corollary 1]).
@@
Let $\cal C$ be a set of categorical random variables defined on a sample space $D$. If $X\colon D\to\Sigma_{X}$ and $Y\colon D\to\Sigma_{Y}$ are two elements in $\cal C$}, we say that they are indiscernible if there is a bijection $h\colon\Sigma_{X}\to\Sigma_{Y}$ such that $Y=h\circ X$ almost everywhere.
@@
Example 3.5 (of indiscernibles).
Consider variable $X$ with values 1, 2 or 3 and variable $Y$ with values A, B or C. Table 5 shows a sample of 10 individuals. Considering the probability given by the observed frequency, we can see that the histogram of $X$ is 0.5, 0.1, 0.4, while the histogram of $Y$ is 0.4, 0.1, 0.5. Here, we can define the bijection $h\colon\{1,2,3\}\to\{A,B,C\}$ given by $h(1)=C;h(2)=B;h(3)=A$, and then the partition defined on the domain by $Y$ and $h\circ X$ are the same.
@@
    Positivity and normalization: $SU(X,Y)\in[0,1]$.
Proof 3.7.
@@
(Condition 1) is straightforward by substitution into equation (1). Likewise, for the reflexivity (Condition 2), using equation (1) we get the self-similarity $SU(X,X)=1\geq 0$.
The self-similarity (Condition 3) establishes that $SU(X,X)$ is greater than or equal to any other value of SU(X, Y). This is indeed true as we have just seen that $SU(X,X)=1$ which is the maximum value achievable by the measure.
Equivalent to the triangle inequality in a distance metric, the triangle inequality (Condition 4) states that the similarity between $X$ and $Z$ through $Y$ is no greater than the direct similarity between $X$ and $Z$ plus the self-similarity of $Y$. Let us start by rewriting the inequalities for the two-variable cases, using entropic ratio notation (see definition in Lemma 2.1). For the variables $X$, $Y$, and $Z$, we get:
@@
b. From here on, use only the left and the middle members. Transpose two terms to the left, getting
@@
(Condition 5) results from applying $SU(X,X)$ = 1 once again, and noting that $X=Y$ is understood as equality in entropy.
Finally, notice that (Condition 6) was already established in Lemma 2.1. This finalizes the proof.
@@
Given a space $\cal C$ of categorical random variables, let $\overline{\cal C}$ be its quotient space defined by indiscernibility. The quantity 1SU(X,Y)1-SU(X,Y) is a normalized distance metric in $\overline{\cal C}$.
Let us notice that, since we are working with categorical variables, it may seem that the metric topology just introduced is discrete. However, this is not the case, as we show in the following theorem: